When a ray of light enters from one medium to another medium it`s velocity is doubled.Find it`s critical angle.

Grade:11

1 Answers

Vikas TU
6 years ago
Dears Student,
As we all know that,
N = 1/sin (critical angle)
And n = c/v
C/v = 1/sin (critical angle)
Critical angle  = v/c
As velocity is get doubled
Critical angle = 2v/c.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)